OpenStudy (anonymous):

y = - 2x + 1... y = - x + 5

OpenStudy (nenadmatematika):

(x,y)=(-4,9)

OpenStudy (anonymous):

You want to use elimination?

OpenStudy (anonymous):

substitution it looks

OpenStudy (anonymous):

Both ways work

OpenStudy (anonymous):

-2x+1=-x+5 -x=4 x=-4

jhonyy9 (jhonyy9):

y=-2x+1 y=-x+5 -2x+1=-x+5 -5+1=-x+2x -4 = x y=-x+5 y=-(-4)+5 y=4+5 y=9 x=-4 y=9